OpenStudy (anonymous):

r-2[1-3(2r+4)]=61 i know the answer but i dont get how to solve it. i have tried like a million different ways

OpenStudy (anonymous):

i did... i am still not understanding...

OpenStudy (kropot72):

r = 3. \[r-2(1-6r-12)=61\] \[r-2+12r+24=61\] 13r = 39 r = 3

OpenStudy (anonymous):

thaNK you so much

OpenStudy (campbell_st):

start with the inner most brackets r - 2[ 1 - 6r - 12] = 61 simplify r - 2[ -6r - 11] = 61 r + 12r + 22 = 61 13r + 22 = 61 now solve